### v1

# Uncertainty quantification in fine-tuned LLMs using LoRA ensembles

Abstract
Fine-tuning large language models can improve task specific performance, although a general understanding of what the fine-tuned model has learned, forgotten and how to trust its predictions is still missing. We derive principled uncertainty quantification for fine-tuned LLMs with posterior approximations using computationally efficient low-rank adaptation ensembles. We analyze three common multiple-choice datasets using low-rank adaptation ensembles based on Mistral-7b, and draw quantitative and qualitative conclusions on their perceived complexity and model efficacy on the different target domains during and after fine-tuning. In particular, backed by the numerical experiments, we hypothesise about signals from entropic uncertainty measures for data domains that are inherently difficult for a given architecture to learn.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) learns conditional distributions of vocabularies useful for generative tasks, text classification and code completion. Trained on a corpus of sequences produced by humans, such as natural language text and program source code, these models have shown remarkable capabilities (Bubeck et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023). As the applications of LLMs continue to explode, informed use of these models hinge on a proper understanding of the uncertainty of their output (Huang et al., 2023; Kuhn et al., 2023; Malinin & Gales, 2021; Ren et al., 2023).
To align an LLM towards particular needs, such as answering factual questions and instructions, a common approach is to fine-tune the model using a specialised dataset targetting human interaction (Ouyang et al., 2022). Fine-tuning involves additional training of a pre-trained LLM on a smaller dataset generated with humans (Zhong et al., 2021) or by other LLMs (Peng et al., 2023; Zhang et al., 2023). By fine-tuning, the model adapts its parameters to better capture the nuances, vocabulary, and style of the target domain(Peng et al., 2023).
Even full-model fine-tuning often requires orders of magnitude less training time, but still incur the same computational complexity and memory usage. Prior studies have shown that pre-trained language models can effectively adapt to specific tasks within smaller parameter subspaces (Li et al., 2018; Aghajanyan et al., 2020), indicating an inherently low rank during the fine-tuning process. To further reduce the training time and computational complexity of fine-tuning, a common method used is Low-Rank Adaptation (LoRA) (Hu et al., 2021). As one member of the family of methods typically referred to as parameter efficient fine-tuning (PEFT), LoRA effectively reduces the number of parameters requiring training by keeping the pre-trained weights unchanged and integrating low-rank trainable matrices into each layer of the LLM transformer architecture.
Key questions arise after fine-tuning: What areas of knowledge remain outside the model’s expertise? What knowledge is retained from the pre-trained model? What knowledge is gained during the fine-tuning process on the target dataset? These questions are fundamental in guiding us towards a more reliable, interpretable and trustworthy application of LLMs.
We contribute to addressing these questions by analyzing the uncertainty in LLMs that have been fine-tuned using LoRA for answering multiple-choice question answers (QAs). These tasks involve definite single-token target labels, which significantly simplifies the analysis while still requiring the model to have a thorough understanding of context. We explore how uncertainty estimates, divided into epistemic model uncertainty and aleatoric data uncertainty, can be utilized to observe changes in the models knowledge regarding the presented data. We use predictive entropy and mutual information to quantify uncertainty, the former containing both aleatoric and epistemic contributions whereas the latter is solely epistemic. These entropic uncertainty measures are calculated for a Bayesian posterior derived from an ensemble of LLMs fine-tuned with LoRA.
We use the pre-trained Mistral-7b (Jiang et al., 2023) model as a prior for the weight distribution, fine-tuning on the CommonsenseQA (CQA) (Talmor et al., 2019), and the Social Sciences as well as STEM components from MMLU (Hendrycks et al., 2021) multiple-choice QA datasets. By quantifying the evolution of knowledge using predictive entropy and mutual information, we show how these measures can be used to reason about complexity of the dataset, and expected model efficacy on the target domain. Code is available (Balabanov & Linander, 2024).
2 Contributions
  • We derive posterior approximations for LLMs fine-tuned on target datasets using ensembles of LoRA members. On the way, we provide a Bayesian interpretation of fine-tuning, early-stopping and conditional generative tasks.
  • We show how uncertainty quantification with entropic measures using the ensemble posteriors can be used to reason about dataset complexity and model efficacy on the target domain.
  • We analyze three common multi-choice QA datasets, CQA, MMLU STEM and MMLU Social Sciences, using LoRA ensemble posteriors derived from a pre-trained Mistral-7b model. The analyzis of the evolution of the entropic uncertainty measures during fine-tuning lets us draw quantitative conclusions on the relative complexity, out-of-distribution behaviour, and model efficacy on the different target domains.
3 Related Work
3.1 Uncertainty quantification in LLMs
There has been significant interest in uncertainty quantification across various tasks and domains in neural networks (Gal & Ghahramani, 2015, 2016; Malinin & Gales, 2018; Ovadia et al., 2019b; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023).
This interest extends to the realm of Large Language Models (LLMs), where accurately quantifying prediction uncertainty is a key focus (Xiao et al., 2022a; Lin et al., 2022; Mielke et al., 2022; Chen & Mueller, 2023; Duan et al., 2023; Huang et al., 2023). The application of LLMs in generative tasks introduces unique challenges, notably in measuring uncertainty of the generative outputs. (Liu et al., 2019; Malinin & Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The disentanglement of uncertainty into aleatoric and epistemic was recently discussed in the context of LLMs (Hou et al., 2023). However, it was done via ensembling of the model inputs rather than the model instances, and not in the context of fine-tuning tasks.
3.2 Fine-tuning in LLMs
Fine-tuning has recently become an integral part of the LLM ecosystem where it is used to target specific tasks such as general instruction-following model, through methods like Reinforcement Learning from Human Feedback (Houlsby et al., 2019; Hu et al., 2021; Liu et al., 2019; Ding et al., 2022, 2023).
The large computational demands of training and fine-tuning LLMs have resulted in development of more efficient techniques, known as parameter-efficient fine-tuning (PEFT)(Liu et al., 2022; Ding et al., 2022, 2023; Shi & Lipani, 2024). These methods typically involve training a small number of additional parameters on top of a fixed, pre-trained LLM, with a prominent approach being the use of low-rank adapters (LoRA) for each weight matrix (Hu et al., 2021).
3.3 LLMs within Bayesian methods
Bayesian inference techniques in neural networks have a long-standing history (Denker et al., 1987; Tishby et al., 1989; Buntine & Weigend, 1991; MacKay, 1991). These methods establish a systematic approach to derive reliable and interpretable uncertainty estimates. Previous research on Bayesian language models has been primarily concentrated on the pretraining phase of language models, rather than on their fine-tuning (Tran et al., 2019; Xue et al., 2021; Cinquin et al., 2021; Zhang et al., 2018; Chen & Li, 2024).
Recent studies have begun exploring the fine-tuning of language models using a Bayesian approach. For instance, in (Fan et al., 2020) and (Zhang et al., 2021), the attention modules are sampled either from simplex-constrained attention distributions or using Bayesian Belief Networks. Neither of these studies address entropic uncertainty measures, nor do they utilize ensembling or PEFT methods for posterior approximation.
(Yang et al., 2024) employs a post-hoc Laplace approximation (Mackay, 1992) to model LoRA parameters for fine-tuning. While this study does use LoRA for fine-tuning, it does not explore entropic uncertainty measures and focuses on the posterior over LoRA parameters rather than the model as a whole, which could limit straightforward Bayesian interpretability. There are also strong indications that deep ensembles provide more accurate posteriors compared to single-model stochastic methods like Laplace and Monte Carlo dropout (Gustafsson et al., 2019; Ovadia et al., 2019a; Fort et al., 2019; Wilson & Izmailov, 2020; Dwaracherla et al., 2022; Balabanov et al., 2023).
3.4 Ensembling LLMs
Two recent studies have explored the use of ensembling in fine-tuning LLMs, specifically focusing on full model fine-tuning method where all weights are optimized (Gleave & Irving, 2022; Sun et al., 2022). This approach has large memory overhead by construction. While uncertainty quantification using variance across the ensembles is considered, their methods lack a Bayesian formalism that could provide a more grounded interpretation and understanding.
An alternative method, BatchEnsemble (Wen et al., 2020), employs a base model with modifications through multiplicative, component-specific rank-1 matrices. This method has been applied to LLMs, but in the context of pre-training rather than fine-tuning (Tran et al., 2022).
There are also recent endeavors using LoRA ensembles for fine-tuning LLMs (Wang et al., 2023; Zhai et al., 2023). Although these studies consider uncertainty quantification, they do so without employing a Bayesian framework and do not clearly separate epistemic and aleatoric components, critial for data interpretability.
4 Bayesian Deep Learning for LLMs
When formulated as a conditional generative model, an LLM can be viewed as a function taking a sequence of tokens $s^{*}$ as input, and predicting a distribution over the set of possible tokens $t^{*}$ from the vocabulary as output. Such a model is typically trained to predict the next token in a sequence.
The prediction should be consistent with our observed data. This is encapsulated in the predictive probability distribution $p(t^{*}|s^{*},\mathcal{D})$, with training data $\mathcal{D}=\{s_{i},t_{i}\}_{i=1}^{N}$, giving a conditional distribution over tokens $t^{*}$. To obtain a specific prediction, we can sample from this distribution, for example, using maximum posterior estimation, $t^{*}=\arg\max_{t}p(t|s^{*},\mathcal{D})$.
Given a model with parameters $\theta$, the predictive distribution $p(t^{*}|s^{*},\mathcal{D})$ can be expressed as follows
[…]
The posterior parameter distribution $p(\theta|\mathcal{D})$ contains information about the epistemic uncertainty associated with model parameters $\theta$, while $p(t^{*}|s^{*},\theta)$ reflects aleatoric uncertainty inherent in the data. Thus, as any data driven model, the output of LLMs has an associated uncertainty stemming from an uncertainty about model parameters, as well as an uncertainty regarding the correct output given the input sequence.
An example of aleatoric uncertainty is given by the input sequence “The color of the sky is ”. One would expect a model trained on a large corpus of text to produce a predictive distribution with support on the colors. This posterior distribution tells us that the model is fairly certain that the next token is a color, but that there is an uncertainty about which color. The uncertainty about the color cannot be decreased by improving the training data, it is inherent in the input sequence.
Moreover, aleatoric uncertainty depends on the architecture of the model. For example, when a current generation LLM is asked a question like “The square root of 123456789 equals ”, we expect the predictive distribution to exhibit high aleatoric uncertainty. Although the answer is definitive, and the model understands the context, it might still be incapable of providing the correct answer. This limitation might be inherent in the models architecture, and cannot be resolved through further training.
On the other hand, the input sequence “Large language models are ” presents a simple conditional completion task, but we would not trust the output predictions from a model trained on a text corpus compiled before the concept was introduced. This uncertainty is not captured in the output distribution, but rather in the uncertainty of the model parameters themselves. This epistemic uncertainty can be systematically calculated within a Bayesian formalism, where it corresponds to shape of the posterior distribution of model parameters conditioned on the training data.
Separating the origin of uncertainty is important to make informed decisions about how to interpret the preditions of LLMs. Should we trust the output for a given sequence at all, and given such trust, what is the inherent uncertainty of the output?
4.1 Fine-tuning
The objective of fine-tuning is to use a specialized dataset to further improve the knowledge of a generally pre-trained model for a given task. This process can be conceptualized as a conditional generative task on a target domain, denoted as $\mathcal{D}_{\text{fine-tuning}}$.
Fine-tuning assumes prior knowledge about the models parameters, expressed as $p(\theta)$. This prior is typically centered around the pre-trained models parameters, $\omega_{\text{pre-trained}}$, to retain the existing knowledge and apply it to the target domain. We assume a normal prior distribution centered on the pre-trained parameter values: $p(\theta)=N(\theta;\omega_{\text{pretrained}},\lambda^{-1}I_{\text{dim}[\theta]})$ with diagonal covariance $I_{\text{dim}[\theta]}$. By adjusting the prior variance $\lambda^{-1}$, we can control the extent to which the model is allowed to deviate from the pre-trained parameter values. In practice, this involves finding a balance between two extremes: a small $\lambda^{-1}$ means that the fine-tuning is largely disregarded, whereas a large $\lambda^{-1}$ implies that the pre-trained models knowledge is not utilized. As with any posterior considerations in a Bayesian formalism, fine-tuning requires careful adjustment of this parameter to target optimal performance for the target tasks.
Note that in this formulation, there is no emphasis on the origin of the pre-trained parameters. Typically, for large language models, the training procedure consists of maximum likelihood optimization combined with supervised reinforcement learning (Touvron et al., 2023b) and the specific details of the training procedure are often not publicly disclosed. In our formulation, we do not need to delve into these origins and instead formally take it as prior knowledge upon which we base our fine-tuning.
4.2 Posterior approximation methods
Variational inference (Barber & Bishop, 1998; Graves, 2011; Blundell et al., 2015) offers a computationally efficient approximation method for the Bayesian posterior. It involves exploring a family of distributions $q_{\omega}(\theta)$, parameterized by $\omega$, to approximate the posterior $p(\theta|\mathcal{D})$. A prevalent approach to identify the optimal $q_{\omega}(\theta)$ in this set involves minimizing the Kullback-Leibler (KL) divergence $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$ (Gawlikowski et al., 2021). This divergence, which measures information loss when $q_{\omega}(\theta)$ approximates $p(\theta|\mathcal{D})$, can be expressed as:
$\begin{split}&\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))=\int\,d\theta\,q_{\omega}(\theta)\,\log\,\frac{q_{\omega}(\theta)}{p(\theta|\mathcal{D})}\\ &=\text{KL}(q_{\omega}(\theta)\,||\,p(\theta))-\mathbb{E}_{q_{\omega}(\theta)}[\log\,p(t|\theta,s)]+C,\\ \end{split}$ (2)
where $\mathcal{D}=\{(t,s)\}$ represents the training data, and $C$ is a constant term independent of $\theta$. Here we applied Bayes’ theorem $p(\theta|t,s)=p(t|\theta,s)\,p(\theta)/p(t|s)$. Minimizing the loss function given in Eq. (2) results in an approximation of the posterior $p(\theta|\mathcal{D})$. This loss function includes two components: (1) the KL divergence against the prior $p(\theta)$ and (2) the expected negative log likelihood (ENLL), a common loss metric for classification tasks.
The optimization problem described in Eq. (2) does not specify the origin of the in-domain (observed) data $\mathcal{D}$, and it remains equally valid for a fine-tuning task when substituting $\mathcal{D}=\mathcal{D}_{\text{fine-tune}}$. However, it’s important to stress that for a fine-tuning task, the prior is assumed to be centered around the pre-trained solution $\omega_{\text{pretrained}}$, whereas for a typical training procedure, the prior is usually considered to be centered around the origin in parameter space.
4.3 Early Stopping
Early stopping can be used to enhance generalization to unseen in-domain data, thereby reducing the effect of overfitting.
In the context of Bayesian deep learning, the objective is to approximate the posterior distribution $p(\theta|\mathcal{D})$, where $\mathcal{D}$ represents the entire task domain. Access to the full domain $\mathcal{D}$ is typically not available, and we work with a subset $\mathcal{D}^{\text{train}}$. Generally, $p(\theta|\mathcal{D}^{\text{train}})$ does not closely mirror $p(\theta|\mathcal{D})$, leading to overfitting, where $p(\theta|\mathcal{D}^{\text{train}})$ only acknowledges the training data and struggles to generalize.
Early stopping improves the approximation of $p(\theta|\mathcal{D})$. Within the variational inference picture, $q_{\omega}(\theta)$ is optimized by minimizing $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}^{\text{train}}))$, but evaluated on the entire task domain $\mathcal{D}$ using $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$. The prior contribution to the KL loss $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$ can be dropped assuming the entire task domain $\mathcal{D}$ to have a large number of samples. In practice, the ENLL contribution to the KL loss is estimated using the validation dataset $\mathcal{D}^{\text{val}}$, assuming it represents the entire task domain effectively and was not part of the training.
5 Deep Ensembles with Low-Rank Adaptation
5.1 Low-Rank Adaptation
Low-Rank Adaptation (LoRA) provides a computationally efficient alternative to full-model fine-tuning by keeping the pre-trained weight matrices static and adding trainable low-rank matrices into each transformer layer, thereby decreasing the total number of parameters requiring training (Hu et al., 2021).
Given a pre-trained weight matrix $W_{\text{pretrained}}$ in the space $\mathbb{R}^{d\times k}$, the adaptations are represented as $W_{\text{pretrained}}+BA$, where $B$ resides in $\mathbb{R}^{d\times r}$, $A$ in $\mathbb{R}^{r\times k}$, and the rank $r$ is substantially smaller than both $d$ and $k$. Throughout this training approach, the variables $A$ and $B$ are tuned, while $W_{\text{pretrained}}$ is kept unchanged.
5.2 LoRA Deep Ensembles
We use LoRA ensembles for approximating the posterior $p(\theta|\mathcal{D}_{\text{fine-tune}})$. Ensembles of LoRA members have been introduced recently (Wang et al., 2023), here we provide a Bayesian treatment.
Deep ensembles can be explicitly interpreted within the Bayesian variational framework by assuming an ansatz in the form of a sum of sharply peaked distributions around parameter realizations $\omega_{k}$, with $k$ indexing ensemble members (Hoffmann & Elster, 2021; Balabanov et al., 2023). In this context, the variational KL loss simplifies to a sum of conventional single-member log likelihood loss terms, augmented with L2 regularization.
We adopt the deep ensemble variational inference formulation to include LoRA reparametrization. The trainable low-rank LoRA matrices $A_{k}$ and $B_{k}$ are not treated as a subset of the variational parameters $\omega_{k}$ of the deep ensemble ansatz. Rather, they are auxiliary degrees of freedom that parameterize these variational parameters. That is, we multiply matrices $A_{k}$ and $B_{k}$ to obtain $W_{k}=(W_{\text{pretrained}}+A_{k}B_{k})\subset\omega_{k}$, and $W_{k}$ is used in the deep ensemble ansatz for approximating the posterior distribution. The LoRA parametrisation helps us to efficiently and inexpensively find good values for $W_{k}\subset\omega_{k}$.
5.3 Prior Selection
The choice of an optimal prior is crucial for obtaining a posterior distribution that exhibits optimal performance. Our approach is to assume a normally distributed prior around the pre-trained model, $p(\theta)=N(\theta;\omega_{\text{pretrained}},\lambda^{-1}I_{\text{dim}[\theta]})$, defined with a certain variance $\lambda^{-1}$. This variance is a hyperparameter that must be carefully chosen based on specific objectives.
An inappropriate selection of the prior can lead to a posterior that performs poorly on the task at hand. Consequently, we adjust the variance of the prior with the aim of optimizing the posterior quality, as measured by the log likelihood on the validation data. Different choice for the priors variance results in adjusting L2 regularization loss according to
$\begin{split}&\text{L2}=\frac{\lambda}{2}\sum_{i}||W^{(i)}-W^{(i)}_{\text{pretrained}}||^{2}=\frac{\lambda}{2}\sum_{i}||B^{(i)}A^{(i)}||^{2},\\ \end{split}$ (3)
where $A^{(i)}$ and $B^{(i)}$ are LoRA matrices associated with the weights $W^{(i)}$.
6 Uncertainty quantification
We quantify total predictive uncertainty, with both aleatoric and epistemic contributions, using predictive entropy
$\begin{split}&H(t^{*}|s^{*},\mathcal{D})=-\sum_{c}p(t^{*}=c|s^{*},\mathcal{D})\log p(t^{*}=c|s^{*},\mathcal{D}),\\ \end{split}$ (4)
where $s^{*}$ is the input string, $\mathcal{D}$ is the training data, and $c$ represents all possible output tokens. The output probabilities $p(c|s^{*},\mathcal{D})$ are computed by averaging softmax outputs from network parameters $\theta$, sampled from the Bayesian posterior $p(\theta|\mathcal{D})$.
We quantify epistemic uncertainty by the average mutual Shannon information between the model parameters and a test data sample, given the training dataset (Mackay, 1992)
$\begin{split}&\text{MI}(\theta,t^{*}|s^{*},\mathcal{D})\\ &=H(\theta|\mathcal{D})-\mathbb{E}_{t^{*}\sim p(t^{*}|s^{*},\mathcal{D})}\left[H(\theta|t^{*},s^{*},\mathcal{D})\right]\\ &=H(t^{*}|s^{*},\mathcal{D})-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}\left[H(t^{*}|s^{*},\theta)\right].\\ \end{split}$ (5)
The mutual information MI in Eq. (5) is the expected change in the posterior parameter distribution when a new data point is added to the training set (first row of RHS) (Houlsby et al., 2011), and can be conveniently calculated in terms of the entropy of the predictive distributions (second row of RHS).
Even though entropic uncertainty measures have been shown to contain valuable information for neural network uncertainty quantification (Linander et al., 2023), one should be aware of unintuitive properties shown in recent studies (Wimmer et al., 2023).
7 Numerical Results
7.1 Datasets
We evaluate our methods on three multiple-choice QA datasets: CommonsenseQA (CQA) (Talmor et al., 2019), and the Social Sciences (SS) as well as STEM components from MMLU (Hendrycks et al., 2021). Detailed information about the training and validation (test) splits is available in Appendix A.
7.2 LoRA Ensembles
We train LoRA ensembles using the pre-trained 7 billion parameter Mistral-7b model (Jiang et al., 2023). The LoRA parametrization was implemented using PEFT (Mangrulkar et al., 2022). Our ensembles contain 5 members $(M=5)$, but we also investigate larger ensembles in our ablation studies. Adhering to the method of Hu et al. (Hu et al., 2021), we applied the adapter $\Delta W=\alpha BA$ exclusively to the query and value matrices within the self-attention modules of Mistral-7b, with $\alpha$ set to 32. The rank was set to $r=8$, which gives $3,407,872$ trainable parameters. The adapter matrices $B$ are initialized to zero, while the entries of $A$ are initialized using a Kaiming-uniform distribution (He et al., 2015).
The ensemble members are trained using the prescribed regularisation from Section 5.3, corresponding to using the pretrained Mistral-7b model as a prior. We use the Adam optimizer, training for 10 epochs with learning step sizes of $5\cdot 10^{-6}$ and $1.25\cdot 10^{-6}$, and batch sizes of 8 and 2 for the CQA and MMLU datasets, respectively. These learning rates are smaller compared to those in (Wang et al., 2023) to better refine the learning curve before the models become over-fitted. The token size limit is fixed at 128 for CQA and 512 for MMLU datasets. We performed an ablation study on L2 regularization $\lambda/2=0.01,0.1,1,10$, finding optimal performance at $\lambda/2=1$ and $10$ for the CQA and MMLU datasets, respectively. The training focuses solely on the single token output representing the answer, with all other token outputs being masked.
To calculate performance metrics and uncertainty estimates, we reduced the output dimension from 32,000 to 6. Five of these dimensions represent QA choice tokens (a, b, c, d, e), and the sixth aggregates the softmax prediction scores associated with all other tokens. We found that, even before fine-tuning, the pretrained Mistral-7b checkpoint assigns a near-zero probability to the sixths class with very few exceptions. This suggests that the model reliably comprehends the format of multiple-choice questions and answers, consistently producing appropriate tokens.
Figure 1: Performance of the LoRA ensembles trained and evaluated on either CQA, MMLU STEM, or MMLU SS dataset. The metrics include accuracy, log likelihood loss, and expected calibration error (ECE). The number of ensemble members is either $M=5$ or $M=1$. For a LoRA ensemble with $M=1$ the average results over 5 distinct realizations are shown.
7.3 Performance metrics
In Fig. 1, we illustrate the performance metrics for LoRA ensembles with both $M=1$ and $M=5$ members. The ensembles are trained on multiple-choice QAs from CQA (first column), MMLU STEM (second column), and MMLU Social Studies datasets (third column).
The ensemble members start over-fitting after a couple of epochs as can be seen by the validation losses in the second row of panels, but interestingly without considerable drop in accuracy in the first row of panels. As will be seen in more detail in Section 7.5, this signals that the ensemble members become more overconfident on predictions that are wrong, corroborated by the increasing expected calibration error with epoch seen in the bottom row of Fig. 1. This is consistent with the recent reports that fine-tuned LLMs often exhibit overconfidence (Jiang et al., 2021; Lin et al., 2022; Xiao et al., 2022b; He et al., 2023; Tian et al., 2023; OpenAI et al., 2023).
For fine-tuning on CQA seen in the first column, the overfitting is significantly reduced for ensembles compared to individual members, as can be seen by the gap between the loss curves for $M=1$ and $M=5$ for later epochs. This means that ensemble members output more confident, but different predictions. This is a common feature of Bayesian posteriors (Blundell et al., 2015; Zhang et al., 2020; Kristiadi et al., 2020; Ober & Aitchison, 2021; Fortuin et al., 2022; Aitchison et al., 2021; Yang et al., 2024). This signals high epistemic uncertainty in this regime, indicating that the validation set is perceived as out-of-domain. There is no significant difference between single member and ensemble metrics for models trained on the MMLU datasets. We attribute this to the fact that the MMLU training datasets are small, see Appendix A, facilitating fast overfitting before any significant generalization is attained.
Figure 2: AUROC computed for the CQA, MMLU STEM, and MMLU Social Studies multiple-choice QA datasets. In each subpanel, one dataset is used for fine-tuning (T) on the training subset and two datasets for the AUROC evaluation (V) on the validation subset. The LoRA ensemble size is $M=5$.
7.4 Uncertainty metrics visualized by AUROC
Area under receiver operator curve (AUROC) is commonly used as a metric for out-of-domain detection (Hendrycks & Gimpel, 2016), but can also be used more generally to measure how well a model can separate two datasets. By evaluating AUROC using uncertainty measures, we can gain understanding of which of the datasets is more consistently perceived by the model as more confusing (using entropy) or as more out-of-domain (using mutual information).
In Fig. 2, we show the AUROC results comparing in- and out-of-domain validation datasets. The ensembles were fine-tuned on the training subset of the CQA, MMLU STEM, and MMLU Social Studies datasets. For the AUROC evaluation, we use the validation subset of these datasets as in-domain samples, and the validation subset of other datasets as out-of-domain.
By analyzing the AUROC curves for the ensemble trained on CQA in the first column of Fig. 2, it is evident that the model consistently perceives samples from CQA as more in-domain compared to those from MMLU datasets, as indicated by an AUROC greater than 0.5. This aligns with expectations. However, the AUROC is not close to 1, suggesting that a significant portion of the MMLU dataset is perceived similarly to CQA in terms of entropy and mutual information, with MMLU Social Sciences appearing closer than MMLU STEM as can be seen by the lower AUROC values. A somewhat similar pattern is observed for the ensemble trained on MMLU Social Sciences.
For the ensemble trained on MMLU STEM in the second column, a peculiar behavior is observed, where the AUROC is below 0.5. We interpret this as follows: MMLU STEM poses greater complexity for the LLM architecture, hence, even with training, it struggles to identify the right features for correct answers. Conversely, the ensemble retains robust knowledge about features pertinent to MMLU SS and CQA. A mutual information score below 0.5 indicates that the ensemble perceives the validation set of MMLU STEM as more out-of-domain compared to the other datasets, even though the model was fine-tuned on the training subset of MMLU STEM.
7.5 Dynamics of uncertainty metrics
To analyze the trends observed through the AUROC curves in more detail, we also produce data-resolved dynamics of uncertainty measures by looking at two-dimensional histogram plots of mutual information and entropy(Linander et al., 2023).
In Fig. 3, we show the epoch evolution of uncertainty measures for a LoRA ensemble that was fine-tuned and evaluated on the CQA dataset. The mutual information (MI) on the vertical axis, and the predictive entropy (Entropy) on the horisontal axis, are binned over every sample of the validation set of CQA and the number of samples in each bin is indicated by the colorbar. We distinguish between incorrect and correct predictions to enhance interpretability. The ensemble attains minimal validation loss at epoch 3, as shown in Fig. 1. Epochs 1 and 6 exhibit underfitting and overfitting, respectively. The true posterior $p(\theta|D)$ should yield nearly zero mutual information when evaluated on the validation dataset, which should be in-domain by design. Samples with high mutual information indicate that parts of the target domain are perceived as out-of-domain by the model, showing the most important regions to consider for improving the training dataset targetting a better posterior approximation. We observe a moderate spread in entropy on the horizontal axis, indicating that the model predicts a broad distribution even though the sample is considered in-domain through their small mutual information. This spread can be attributed to the questions where the model does not have enough context or expressive power for a distinct answer.
Fig. 4 shows the uncertainty histograms for MMLU STEM and MMLU SS, calculated using an ensemble trained on CQA. Starting with MMLU STEM in the first and second row of panels, we draw two conclusions. First, out of the wrongly classified samples for the optimal epoch 3 shown in the second row, second column, there is a large fraction with high predictive entropy and low mutual information compared to the more moderate spread of predictive entropy in the corresponding panel for MMLU SS in the last row, second column. This affirms conclusions from Section 7.4 regarding the relatively higher complexity of STEM compared to SS. The ensemble trained on CQA perceives the STEM samples equally in-domain as the SS samples, as evident by the similar mutual information, but predict a broader distribution indicating a lack of knowledge. Secondly, in the over-fitting regime corresponding to the last column, the mutual information of the wrongly classified samples of the second row increase more than the correctly classified samples in the first row. This indicates that the ensemble forgets the samples with high predictive entropy quicker, i.e. it forgets the confusing samples faster.
Figure 3: Histograms of predictive entropy and mutual information for a LoRA ensemble trained and evaluated on the CQA dataset. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. This figure illustrates the evolution of uncertainty measures for in-domain data across training epochs, differentiated by correct and incorrect predictions. We also depict the corresponding mean (red) and median (green) entropy and mutual information values.
Figure 4: Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA and evaluated on MMLU STEM and MMLU SS datasets. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. This panels illustrates the evolution (left to right) of uncertainty measures for out-of-training-domain data across training epochs, differentiated by correct and incorrect predictions. We also depict the corresponding mean (red) and median (green) entropy and mutual information values.
Continuing with MMLU SS in the third and fourth row, we observe a decrease in predictive entropy in the over-fitting regime, indicating that the ensemble does find features giving confident but wrong predictions.
Interestingly, the evolution of entropic quantities shows a quantitative difference: For MMLU STEM, there is a significant density of high-entropy, low-mutual-information samples at epoch 6, whereas for MMLU SS, the data with high entropy is more dispersed in terms of mutual information. This suggests that the fine-tuned model perceives confusing samples from MMLU STEM and MMLU SS differently, forgetting the confusing samples from MMLU SS more rapidly.
8 Conclusions
We derived a principled, and practically efficient method for posterior approximation of fine-tuned LLMs using ensembles of low-rank adaptation models. By analysing the evolution of entropic uncertainty measures during fine-tuning, we identified signals of dataset complexity and signs of architecture limitations.
The methods presented are generally applicable to different fine-tuning scenarios, and provide a systematic treatment of prediction uncertainty for fine-tuned LLMs.
9 Acknowledgements
HL was supported by a grant from the KAW Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at C3SE partially funded by the Swedish Research Council through grant agreements no. 2022-06725 and no. 2018-05973. OB would like to thank Ihor Balabanov for valuable discussions.
Appendix A Datasets
A.1 Size
For MMLU STEM and Social Studies datasets, we merge the development and validation sets to create the training set and use the original test split as the validation set, as in (Wang et al., 2023). For CQA dataset we use the standard training and validation splits. The topics in the MMLU dataset are categorized into STEM, Social Studies, Humanities, and Other, following the classification defined in (Hendrycks et al., 2021).
Dataset Train Size Validation Size
CQA 9741 1221
MMLU SS 397 3077
MMLU STEM 430 3153
Table 1: Summary of the CQA and MMLU dataset sizes.
A.2 Examples
The datasets have been preprocessed into a single, consistently formatted prompt question-answer. Below, we present examples of these questions for each of the considered dataset.
Q: A revolving door is convenient for two direction travel, but it also serves as a security measure at a what? Answer Choices: (a) bank (b) library (c) department store (d) mall (e) new york A: (a). CommonsenseQA
Q: Which one of the following is the most appropriate definition of a 99% confidence interval? Answer Choices: (a) 99% of the time in repeated samples, the interval would contain the true value of the parameter (b) 99% of the time in repeated samples, the interval would contain the estimated value of the parameter (c) 99% of the time in repeated samples, the null hypothesis will be rejected (d) 99% of the time in repeated samples, the null hypothesis will not be rejected when it was false A: (a). MMLU SS
Q: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field. Answer Choices: (a) 0 (b) 1 (c) 2 (d) 3 A: (b). MMLU STEM

### v2

@@
 We derive principled uncertainty quantification for fine-tuned LLMs with posterior approximations using computationally efficient low-rank adaptation ensembles.
-We analyze three common multiple-choice datasets using low-rank adaptation ensembles based on Mistral-7b, and draw quantitative and qualitative conclusions on their perceived complexity and model efficacy on the different target domains during and after fine-tuning.
-In particular, backed by the numerical experiments, we hypothesise about signals from entropic uncertainty measures for data domains that are inherently difficult for a given architecture to learn.
-Machine Learning, ICML
+We analyze three common multiple-choice datasets using low-rank adaptation ensembles based on Mistral-7b, and draw quantitative and qualitative conclusions on their perceived complexity and balance between retained prior knowledge and domain specific adaptation during and after fine-tuning.
+We identify unexpected retention of acquired knowledge during fine-tuning in the overfitting regime.
+Code: github.com/oleksandr-balabanov/equivariant-posteriors
 1 Introduction
 Large language models (LLMs) learns conditional distributions of vocabularies useful for generative tasks, text classification and code completion.
-Trained on a corpus of sequences produced by humans, such as natural language text and program source code, these models have shown remarkable capabilities (Bubeck et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023).
+Trained on a corpus of sequences such as natural language text and program source code, these models have shown remarkable capabilities (Bubeck et al., 2023; Touvron et al., 2023a; b; Jiang et al., 2023).
 As the applications of LLMs continue to explode, informed use of these models hinge on a proper understanding of the uncertainty of their output (Huang et al., 2023; Kuhn et al., 2023; Malinin & Gales, 2021; Ren et al., 2023).
 To align an LLM towards particular needs, such as answering factual questions and instructions, a common approach is to fine-tune the model using a specialised dataset targetting human interaction (Ouyang et al., 2022).
-Fine-tuning involves additional training of a pre-trained LLM on a smaller dataset generated with humans (Zhong et al., 2021) or by other LLMs (Peng et al., 2023; Zhang et al., 2023).
+Fine-tuning involves additional training of a pre-trained LLM on a smaller dataset generated with humans (Zhong et al., 2021) or by other LLMs (Peng et al., 2023; Zhang et al., 2023a).
 By fine-tuning, the model adapts its parameters to better capture the nuances, vocabulary, and style of the target domain(Peng et al., 2023).
 Even full-model fine-tuning often requires orders of magnitude less training time, but still incur the same computational complexity and memory usage.
-Prior studies have shown that pre-trained language models can effectively adapt to specific tasks within smaller parameter subspaces (Li et al., 2018; Aghajanyan et al., 2020), indicating an inherently low rank during the fine-tuning process.
+Prior studies have shown that pre-trained language models can effectively adapt to specific tasks within smaller parameter subspaces Li et al.
+(2018); Aghajanyan et al.
+(2020), indicating an inherently low rank during the fine-tuning process.
 To further reduce the training time and computational complexity of fine-tuning, a common method used is Low-Rank Adaptation (LoRA) (Hu et al., 2021).
@@
 These tasks involve definite single-token target labels, which significantly simplifies the analysis while still requiring the model to have a thorough understanding of context.
-We explore how uncertainty estimates, divided into epistemic model uncertainty and aleatoric data uncertainty, can be utilized to observe changes in the models knowledge regarding the presented data.
-We use predictive entropy and mutual information to quantify uncertainty, the former containing both aleatoric and epistemic contributions whereas the latter is solely epistemic.
+We use predictive entropy and mutual information to quantify inherent- and model uncertainty.
 These entropic uncertainty measures are calculated for a Bayesian posterior derived from an ensemble of LLMs fine-tuned with LoRA.
-We use the pre-trained Mistral-7b (Jiang et al., 2023) model as a prior for the weight distribution, fine-tuning on the CommonsenseQA (CQA) (Talmor et al., 2019), and the Social Sciences as well as STEM components from MMLU (Hendrycks et al., 2021) multiple-choice QA datasets.
-By quantifying the evolution of knowledge using predictive entropy and mutual information, we show how these measures can be used to reason about complexity of the dataset, and expected model efficacy on the target domain.
-Code is available (Balabanov & Linander, 2024).
+Using the pre-trained Mistral-7b Jiang et al.
+(2023) model as a prior, we fine-tune on the CommonsenseQA (CQA) Talmor et al.
+(2019) and evaluate its performance on both CQA and the Social Sciences and STEM components of the MMLU Hendrycks et al.
+(2021) multiple-choice QA datasets.
+By quantifying the evolution of knowledge using predictive entropy and mutual information, we demonstrate how these measures can be used to reason about the balance between retained prior knowledge and acquired domain specific knowledge during and after fine-tuning.
 2 Contributions
+  • We show how uncertainty quantification with entropic measures can be used to reason about the balance between retained knowledge and domain specific adaptation.
   • We derive posterior approximations for LLMs fine-tuned on target datasets using ensembles of LoRA members.
 On the way, we provide a Bayesian interpretation of fine-tuning, early-stopping and conditional generative tasks.
-  • We show how uncertainty quantification with entropic measures using the ensemble posteriors can be used to reason about dataset complexity and model efficacy on the target domain.
-  • We analyze three common multi-choice QA datasets, CQA, MMLU STEM and MMLU Social Sciences, using LoRA ensemble posteriors derived from a pre-trained Mistral-7b model.
-The analyzis of the evolution of the entropic uncertainty measures during fine-tuning lets us draw quantitative conclusions on the relative complexity, out-of-distribution behaviour, and model efficacy on the different target domains.
+  • We analyze three common multi-choice QA datasets, CQA, MMLU STEM and MMLU Social Sciences, using LoRA ensemble posteriors derived from a pre-trained Mistral-7b model, and identify unexpected retention of acquired knowledge in the overfitting regime.
 3 Related Work
 3.1 Uncertainty quantification in LLMs
-There has been significant interest in uncertainty quantification across various tasks and domains in neural networks (Gal & Ghahramani, 2015, 2016; Malinin & Gales, 2018; Ovadia et al., 2019b; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023).
-This interest extends to the realm of Large Language Models (LLMs), where accurately quantifying prediction uncertainty is a key focus (Xiao et al., 2022a; Lin et al., 2022; Mielke et al., 2022; Chen & Mueller, 2023; Duan et al., 2023; Huang et al., 2023).
+There has been significant interest in uncertainty quantification across various tasks and domains in neural networks Gal & Ghahramani (2015; 2016); Malinin & Gales (2018); Ovadia et al.
+(2019b); Malinin et al.
+(2021); Lin et al.
+(2022); Kuhn et al.
+(2023); Lin et al.
+(2023).
+This interest extends to the realm of Large Language Models (LLMs), where accurately quantifying prediction uncertainty is a key focus Xiao et al.
+(2022a); Lin et al.
+(2022); Mielke et al.
+(2022); Chen & Mueller (2023); Duan et al.
+(2023); Huang et al.
+(2023); Shorinwa et al.
+(2024).
 The application of LLMs in generative tasks introduces unique challenges, notably in measuring uncertainty of the generative outputs.
-(Liu et al., 2019; Malinin & Gales, 2021; Kuhn et al., 2023; Lin et al., 2023).
-The disentanglement of uncertainty into aleatoric and epistemic was recently discussed in the context of LLMs (Hou et al., 2023).
+Liu et al.
+(2019); Malinin & Gales (2021); Kuhn et al.
+(2023); Lin et al.
+(2023).
+The disentanglement of uncertainty into aleatoric and epistemic was recently discussed in the context of LLMs Hou et al.
+(2023); Ling et al.
+(2024).
 However, it was done via ensembling of the model inputs rather than the model instances, and not in the context of fine-tuning tasks.
 3.2 Fine-tuning in LLMs
-Fine-tuning has recently become an integral part of the LLM ecosystem where it is used to target specific tasks such as general instruction-following model, through methods like Reinforcement Learning from Human Feedback (Houlsby et al., 2019; Hu et al., 2021; Liu et al., 2019; Ding et al., 2022, 2023).
-The large computational demands of training and fine-tuning LLMs have resulted in development of more efficient techniques, known as parameter-efficient fine-tuning (PEFT)(Liu et al., 2022; Ding et al., 2022, 2023; Shi & Lipani, 2024).
-These methods typically involve training a small number of additional parameters on top of a fixed, pre-trained LLM, with a prominent approach being the use of low-rank adapters (LoRA) for each weight matrix (Hu et al., 2021).
+Fine-tuning has recently become an integral part of the LLM ecosystem where it is used to target specific tasks such as general instruction-following model, through methods like Reinforcement Learning from Human Feedback Houlsby et al.
+(2019); Hu et al.
+(2021); Liu et al.
+(2019); Ding et al.
+(2022; 2023).
+The large computational demands of training and fine-tuning LLMs have resulted in development of more efficient techniques, known as parameter-efficient fine-tuning (PEFT)Liu et al.
+(2022); Ding et al.
+(2022; 2023); Shi & Lipani (2024).
+These methods typically involve training a small number of additional parameters on top of a fixed, pre-trained LLM, with a prominent approach being the use of low-rank adapters (LoRA) for each weight matrix Hu et al.
+(2021).
 3.3 LLMs within Bayesian methods
-Bayesian inference techniques in neural networks have a long-standing history (Denker et al., 1987; Tishby et al., 1989; Buntine & Weigend, 1991; MacKay, 1991).
+Bayesian inference techniques in neural networks have a long-standing history Denker et al.
+(1987); Tishby et al.
+(1989); Buntine & Weigend (1991); MacKay (1991).
 These methods establish a systematic approach to derive reliable and interpretable uncertainty estimates.
-Previous research on Bayesian language models has been primarily concentrated on the pretraining phase of language models, rather than on their fine-tuning (Tran et al., 2019; Xue et al., 2021; Cinquin et al., 2021; Zhang et al., 2018; Chen & Li, 2024).
+Previous research on Bayesian language models has been primarily concentrated on the pretraining phase of language models, rather than on their fine-tuning Tran et al.
+(2019); Xue et al.
+(2021); Cinquin et al.
+(2021); Zhang et al.
+(2018); Chen & Li (2024).
 Recent studies have begun exploring the fine-tuning of language models using a Bayesian approach.
-For instance, in (Fan et al., 2020) and (Zhang et al., 2021), the attention modules are sampled either from simplex-constrained attention distributions or using Bayesian Belief Networks.
+For instance, in Fan et al.
+(2020) and Zhang et al.
+(2021), the attention modules are sampled either from simplex-constrained attention distributions or using Bayesian Belief Networks.
 Neither of these studies address entropic uncertainty measures, nor do they utilize ensembling or PEFT methods for posterior approximation.
-(Yang et al., 2024) employs a post-hoc Laplace approximation (Mackay, 1992) to model LoRA parameters for fine-tuning.
+Yang et al.
+(2024) employs a post-hoc Laplace approximation Mackay (1992) to model LoRA parameters for fine-tuning.
 While this study does use LoRA for fine-tuning, it does not explore entropic uncertainty measures and focuses on the posterior over LoRA parameters rather than the model as a whole, which could limit straightforward Bayesian interpretability.
-There are also strong indications that deep ensembles provide more accurate posteriors compared to single-model stochastic methods like Laplace and Monte Carlo dropout (Gustafsson et al., 2019; Ovadia et al., 2019a; Fort et al., 2019; Wilson & Izmailov, 2020; Dwaracherla et al., 2022; Balabanov et al., 2023).
+There are also strong indications that deep ensembles provide more accurate posteriors compared to single-model stochastic methods like Laplace and Monte Carlo dropout Gustafsson et al.
+(2019); Ovadia et al.
+(2019a); Fort et al.
+(2019b); Wilson & Izmailov (2020); Dwaracherla et al.
+(2022); Balabanov et al.
+(2023).
+For a more detailed discussion, see Sec.
+5.4.
 3.4 Ensembling LLMs
-Two recent studies have explored the use of ensembling in fine-tuning LLMs, specifically focusing on full model fine-tuning method where all weights are optimized (Gleave & Irving, 2022; Sun et al., 2022).
+Two recent studies have explored the use of ensembling in fine-tuning LLMs, specifically focusing on full model fine-tuning method where all weights are optimized Gleave & Irving (2022); Sun et al.
+(2022).
 This approach has large memory overhead by construction.
 While uncertainty quantification using variance across the ensembles is considered, their methods lack a Bayesian formalism that could provide a more grounded interpretation and understanding.
-An alternative method, BatchEnsemble (Wen et al., 2020), employs a base model with modifications through multiplicative, component-specific rank-1 matrices.
-This method has been applied to LLMs, but in the context of pre-training rather than fine-tuning (Tran et al., 2022).
-There are also recent endeavors using LoRA ensembles for fine-tuning LLMs (Wang et al., 2023; Zhai et al., 2023).
-Although these studies consider uncertainty quantification, they do so without employing a Bayesian framework and do not clearly separate epistemic and aleatoric components, critial for data interpretability.
+An alternative method, BatchEnsemble Wen et al.
+(2020), employs a base model with modifications through multiplicative, component-specific rank-1 matrices.
+This method has been applied to LLMs, but in the context of pre-training rather than fine-tuning Tran et al.
+(2022).
+There are also recent endeavors employing LoRA ensembles for fine-tuning LLMs Wang et al.
+(2023); Zhai et al.
+(2023).
+Although these studies address uncertainty quantification, they do so without adopting a Bayesian framework and do not clearly distinguish between epistemic and aleatoric components, which is critical for data interpretability.
+In another work Arteaga et al.
+(2024), a separate predictor is trained on the LoRA ensembles uncertainty metrics to detect hallucinations.
+Niu et al.
+(2024) introduces a mixture-of-expert framework (UQ4CT) to explicitly capture and calibrate model uncertainty during fine-tuning.
+All these works are highly relevant yet differ from ours, which provides a principled Bayesian treatment for LoRA ensembles during fine-tuning and examines the evolution of uncertainty metrics to draw conclusions about the balance between retained prior knowledge and newly acquired domain adaptation.
 4 Bayesian Deep Learning for LLMs
@@
 To obtain a specific prediction, we can sample from this distribution, for example, using maximum posterior estimation, $t^{*}=\arg\max_{t}p(t|s^{*},\mathcal{D})$.
-Given a model with parameters $\theta$, the predictive distribution $p(t^{*}|s^{*},\mathcal{D})$ can be expressed as follows
+Given a model with parameters $\theta$, the predictive distribution
 […]
-The posterior parameter distribution $p(\theta|\mathcal{D})$ contains information about the epistemic uncertainty associated with model parameters $\theta$, while $p(t^{*}|s^{*},\theta)$ reflects aleatoric uncertainty inherent in the data.
-Thus, as any data driven model, the output of LLMs has an associated uncertainty stemming from an uncertainty about model parameters, as well as an uncertainty regarding the correct output given the input sequence.
-An example of aleatoric uncertainty is given by the input sequence “The color of the sky is ”.
-One would expect a model trained on a large corpus of text to produce a predictive distribution with support on the colors.
-This posterior distribution tells us that the model is fairly certain that the next token is a color, but that there is an uncertainty about which color.
-The uncertainty about the color cannot be decreased by improving the training data, it is inherent in the input sequence.
-Moreover, aleatoric uncertainty depends on the architecture of the model.
-For example, when a current generation LLM is asked a question like “The square root of 123456789 equals ”, we expect the predictive distribution to exhibit high aleatoric uncertainty.
-Although the answer is definitive, and the model understands the context, it might still be incapable of providing the correct answer.
-This limitation might be inherent in the models architecture, and cannot be resolved through further training.
-On the other hand, the input sequence “Large language models are ” presents a simple conditional completion task, but we would not trust the output predictions from a model trained on a text corpus compiled before the concept was introduced.
-This uncertainty is not captured in the output distribution, but rather in the uncertainty of the model parameters themselves.
-This epistemic uncertainty can be systematically calculated within a Bayesian formalism, where it corresponds to shape of the posterior distribution of model parameters conditioned on the training data.
-Separating the origin of uncertainty is important to make informed decisions about how to interpret the preditions of LLMs.
-Should we trust the output for a given sequence at all, and given such trust, what is the inherent uncertainty of the output?
+is obtained by marginalizing over the posterior distribution of the parameters.
+The posterior $p(\theta\mid\mathcal{D})$ represents the uncertainty associated with model parameters $\theta$, quantifying how likely each choice of $\theta$ is, given the observed data.
+For a fixed $\theta$, the term $p(t^{*}\mid s^{*},\theta)$ captures data uncertainty, reflecting the variability in the model’s output $t^{*}$ for the input $s^{*}$.
+By integrating over all possible parameter values $\theta$, the final predictive distribution combines both uncertainty in the parameters with the inherent variability in the data Gawlikowski et al.
+(2023).
 4.1 Fine-tuning
@@
 Note that in this formulation, there is no emphasis on the origin of the pre-trained parameters.
-Typically, for large language models, the training procedure consists of maximum likelihood optimization combined with supervised reinforcement learning (Touvron et al., 2023b) and the specific details of the training procedure are often not publicly disclosed.
+Typically, for large language models, the training procedure consists of maximum likelihood optimization combined with supervised reinforcement learning Touvron et al.
+(2023b) and the specific details of the training procedure are often not publicly disclosed.
 In our formulation, we do not need to delve into these origins and instead formally take it as prior knowledge upon which we base our fine-tuning.
 4.2 Posterior approximation methods
-Variational inference (Barber & Bishop, 1998; Graves, 2011; Blundell et al., 2015) offers a computationally efficient approximation method for the Bayesian posterior.
+Variational inference Barber & Bishop (1998); Graves (2011); Blundell et al.
+(2015) offers a computationally efficient approximation method for the Bayesian posterior.
 It involves exploring a family of distributions $q_{\omega}(\theta)$, parameterized by $\omega$, to approximate the posterior $p(\theta|\mathcal{D})$.
-A prevalent approach to identify the optimal $q_{\omega}(\theta)$ in this set involves minimizing the Kullback-Leibler (KL) divergence $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$ (Gawlikowski et al., 2021).
+A prevalent approach to identify the optimal $q_{\omega}(\theta)$ in this set involves minimizing the Kullback-Leibler (KL) divergence $\text{KL}(q_{\omega}(\theta)\,||\,p(\theta|\mathcal{D}))$ Gawlikowski et al.
+(2023).
 This divergence, which measures information loss when $q_{\omega}(\theta)$ approximates $p(\theta|\mathcal{D})$, can be expressed as:
@@
 5.1 Low-Rank Adaptation
-Low-Rank Adaptation (LoRA) provides a computationally efficient alternative to full-model fine-tuning by keeping the pre-trained weight matrices static and adding trainable low-rank matrices into each transformer layer, thereby decreasing the total number of parameters requiring training (Hu et al., 2021).
+Low-Rank Adaptation (LoRA) provides a computationally efficient alternative to full-model fine-tuning by keeping the pre-trained weight matrices static and adding trainable low-rank matrices into each transformer layer, thereby decreasing the total number of parameters requiring training Hu et al.
+(2021).
 Given a pre-trained weight matrix $W_{\text{pretrained}}$ in the space $\mathbb{R}^{d\times k}$, the adaptations are represented as $W_{\text{pretrained}}+BA$, where $B$ resides in $\mathbb{R}^{d\times r}$, $A$ in $\mathbb{R}^{r\times k}$, and the rank $r$ is substantially smaller than both $d$ and $k$.
@@
 5.2 LoRA Deep Ensembles
-We use LoRA ensembles for approximating the posterior $p(\theta|\mathcal{D}_{\text{fine-tune}})$.
-Ensembles of LoRA members have been introduced recently (Wang et al., 2023), here we provide a Bayesian treatment.
-Deep ensembles can be explicitly interpreted within the Bayesian variational framework by assuming an ansatz in the form of a sum of sharply peaked distributions around parameter realizations $\omega_{k}$, with $k$ indexing ensemble members (Hoffmann & Elster, 2021; Balabanov et al., 2023).
-In this context, the variational KL loss simplifies to a sum of conventional single-member log likelihood loss terms, augmented with L2 regularization.
-We adopt the deep ensemble variational inference formulation to include LoRA reparametrization.
+We use LoRA ensembles to approximate the posterior $p(\theta|\mathcal{D}_{\text{fine-tune}})$.
+Ensembles of LoRA members have been introduced recently Wang et al.
+(2023), here we provide a Bayesian treatment.
+Deep ensembles can be explicitly interpreted in a Bayesian variational framework by assuming an ansatz consisting of a mixture of sharply peaked Gaussians centered at each parameter realization $\omega_{k}$, with $k$ indexing ensemble members Hoffmann & Elster (2021); Balabanov et al.
+(2023).
+With this ansatz, the variational objective reduces to the sum of each member’s log-likelihood term plus an L2 regularization.
+For additional context, deep ensembles can alternatively be viewed as a particle-based variational inference scheme, where each member evolves under Wasserstein gradient descent to minimize the KL divergence D' Angelo & Fortuin (2021); Wild et al.
+(2023).
+Both interpretations naturally give rise to a repulsive contribution, stemming from the prior term in the KL divergence, that encourages diversity and improves the posterior approximation Balabanov et al.
+(2023); D' Angelo & Fortuin (2021); Wild et al.
+(2023).
+In the present work, we omit this repulsive term and defer its exploration to future research.
+We build on the variational deep‐ensemble framework of Hoffmann et al.Hoffmann & Elster (2021) and Balabanov et al.Balabanov et al.
+(2023), extending it to incorporate the LoRA reparametrization.
 The trainable low-rank LoRA matrices $A_{k}$ and $B_{k}$ are not treated as a subset of the variational parameters $\omega_{k}$ of the deep ensemble ansatz.
@@
 Consequently, we adjust the variance of the prior with the aim of optimizing the posterior quality, as measured by the log likelihood on the validation data.
-Different choice for the priors variance results in adjusting L2 regularization loss according to
+Different choice for the prior variance results in adjusting L2 regularization loss Hoffmann & Elster (2021); Balabanov et al.
+(2023).
+In the case of LoRA ensembles, this loss is given by
 $\begin{split}&\text{L2}=\frac{\lambda}{2}\sum_{i}||W^{(i)}-W^{(i)}_{\text{pretrained}}||^{2}=\frac{\lambda}{2}\sum_{i}||B^{(i)}A^{(i)}||^{2},\\ \end{split}$ (3)
 where $A^{(i)}$ and $B^{(i)}$ are LoRA matrices associated with the weights $W^{(i)}$.
+5.4 Posterior Quality of LoRA Ensembles
+LoRA deep ensembles approximate the true posterior over model parameters for the fine-tuning task via variational inference, and the trustworthiness of the resulting uncertainty estimates depends on how closely this approximation matches $p(\theta|\mathcal{D}_{\text{fine-tune}})$.
+Hamiltonian Monte Carlo (HMC) remains the gold standard for posterior estimation because it generates nearly exact samples and thus yields the most faithful uncertainty estimates.
+In principle, one would evaluate any approximation by measuring its agreement and variance relative to the HMC posterior.
+In practice, HMC-derived posteriors do not always correlate with the best predictive performance: they can be poorly calibrated and sensitive to distributional shifts Izmailov et al.
+(2021), and stochastic ensembles can outperform them in accuracy and NLL Balabanov et al.
+(2023).
+This does not diminish their status as the best available posterior approximation, and as such they are the target for any potential approximation method.
+This closeness can be quantified using agreement and variance metrics Izmailov et al.
+(2021).
+Unfortunately, running HMC on large language models and datasets is prohibitively expensive, placing it beyond our computational budget and preventing direct comparison.
+Although direct comparison is not possible, prior work has shown that multimodal methods such as deep ensembles yield accurate posterior approximations and significantly outperform single-mode approaches Fort et al.
+(2019a); Izmailov et al.
+(2021); Balabanov et al.
+(2023).
+Building on these insights, we argue that our LoRA ensemble–based method delivers high-quality posterior estimates under realistic training and inference constraints, while also surpassing single-mode baselines such as the Laplace approximation, MC dropout, and related techniques.
+We acknowledge that Yang et al.
+Yang et al.
+(2024) examined LoRA-tuned LLMs primarily through a single-mode Laplace approximation, comparing it against several other methods—including a small ensemble—and demonstrated its performance benefits.
+Although their results are pertinent to performance-centric applications, they do not address the uncertainty quantification focus of our study.
+In fact, the Laplace method’s apparent robustness to distribution shifts and “good” calibration may actually indicate a poor approximation of the true posterior Izmailov et al.
+(2021).
 6 Uncertainty quantification
-We quantify total predictive uncertainty, with both aleatoric and epistemic contributions, using predictive entropy
+6.1 Epistemic model and aleatoric data uncertainty
+Uncertainty quantification is an important topic that assesses the reliability of model predictions (Gal & Ghahramani, 2016).
+In general, the total uncertainty in a model’s prediction can be attributed to multiple sources.
+One widely used categorization is the decomposition into epistemic (or model) uncertainty—arising from insufficient knowledge or data to fully learn the underlying distribution—and aleatoric (or data) uncertainty—stemming from the inherent noise and randomness in the data generation process.
+An example of aleatoric data uncertainty is given by the input sequence “There is a single item in the box.
+The color of that item is ”.
+One would expect a model trained on a large corpus of text to produce a predictive distribution with support on the different colors.
+This posterior distribution tells us that the model is fairly certain that the next token is a color, but that there is an uncertainty about which color.
+The uncertainty about the color cannot be decreased by improving the training data; it is inherent in the input sequence.
+On the other hand, the input sequence “Large language models are ” presents a simple conditional completion task, but we would not trust the output predictions from a model trained on a text corpus compiled before the concept was introduced.
+This epistemic model uncertainty is not captured in the output distribution, but rather in the uncertainty of the model parameters themselves.
+This model uncertainty can be systematically calculated within a Bayesian formalism, where it corresponds to the shape of the posterior distribution of model parameters conditioned on the training data.
+6.2 Entropic uncertainty metrics
+Depending on the approach used to estimate uncertainty, existing methods can be broadly divided into intrinsic and extrinsic methods.
+Extrinsic methods involve post-hoc analysis with auxiliary models to estimate uncertainty Lahlou et al.
+(2023); Kristiadi et al.
+(2021).
+In contrast, intrinsic methods incorporate uncertainty estimation directly within the model.
+Here, we employ a common intrinsic methodology based on entropic metrics Malinin & Gales (2021) to evaluate uncertainty within a Bayesian formulation, applying it to our LoRA ensemble approximation of the Bayesian posterior for fine-tuned models.
+The total predictive uncertainty, encompassing both aleatoric data and epistemic model components, can be quantified using the total predictive entropy:
 $\begin{split}&H(t^{*}|s^{*},\mathcal{D})=-\sum_{c}p(t^{*}=c|s^{*},\mathcal{D})\log p(t^{*}=c|s^{*},\mathcal{D}),\\ \end{split}$ (4)
-where $s^{*}$ is the input string, $\mathcal{D}$ is the training data, and $c$ represents all possible output tokens.
-The output probabilities $p(c|s^{*},\mathcal{D})$ are computed by averaging softmax outputs from network parameters $\theta$, sampled from the Bayesian posterior $p(\theta|\mathcal{D})$.
-We quantify epistemic uncertainty by the average mutual Shannon information between the model parameters and a test data sample, given the training dataset (Mackay, 1992)
-$\begin{split}&\text{MI}(\theta,t^{*}|s^{*},\mathcal{D})\\ &=H(\theta|\mathcal{D})-\mathbb{E}_{t^{*}\sim p(t^{*}|s^{*},\mathcal{D})}\left[H(\theta|t^{*},s^{*},\mathcal{D})\right]\\ &=H(t^{*}|s^{*},\mathcal{D})-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}\left[H(t^{*}|s^{*},\theta)\right].\\ \end{split}$ (5)
-The mutual information MI in Eq.
-(5) is the expected change in the posterior parameter distribution when a new data point is added to the training set (first row of RHS) (Houlsby et al., 2011), and can be conveniently calculated in terms of the entropy of the predictive distributions (second row of RHS).
-Even though entropic uncertainty measures have been shown to contain valuable information for neural network uncertainty quantification (Linander et al., 2023), one should be aware of unintuitive properties shown in recent studies (Wimmer et al., 2023).
+where $s^{*}$ is the input string, $\mathcal{D}$ is the training dataset, and $c$ denotes all possible output tokens.
+The output probabilities $p(c\mid s^{*},\mathcal{D})$ are computed by averaging softmax outputs across model parameters $\theta$ sampled from the Bayesian posterior $p(\theta\mid\mathcal{D})$.
+Epistemic model uncertainty is defined as the variability in predictions that arises from sampling model parameters $\theta$ from the posterior distribution $p(\theta\mid\mathcal{D})$.
+This variability captures how the distribution of predicted outputs shifts with each parameter set, weighted by its posterior probability.
+A common metric for this purpose is the average mutual information between the model parameters and output, conditioned on the training data Mackay (1992); Malinin & Gales (2021):
+$\begin{split}&\text{MI}(\theta,t^{*}|s^{*},\mathcal{D})\\ &=H(t^{*}|s^{*},\mathcal{D})-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}\left[H(t^{*}|s^{*},\theta)\right].\\ \end{split}$ (5)
+For ensemble posteriors, zero mutual information implies that all ensemble members yield the same predictive distribution, indicating no variability in the outputs and thus zero epistemic uncertainty.
+On the contrary, if the mutual information is high, there is significant variability in the predictive distribution across the ensemble, reflecting high epistemic uncertainty.
+Using these metrics, we can identify three regimes that describe how the model perceives data:
+  • Known-certain: Low total and epistemic uncertainty; the posterior ensemble strongly agrees on a narrow, confident prediction distribution.
+  • Known-uncertain: High total uncertainty, low epistemic uncertainty; uncertainty arises from data noise or model expressiveness limitations, with the ensemble members agreeing on a broad distribution.
+  • Unknown: High total and epistemic uncertainty; the posterior members disagree on the prediction distribution, highlighting a lack of knowledge.
+In this work, we use uncertainty quantification from entropic metrics of the LoRA ensemble posterior approximation to better understand how fine-tuning reshapes the knowledge of the model, revealing shifts in data characteristics and how the model adapts to new information.
+In particular, we study the dynamics of the uncertainty metrics during fine-tuning.
 7 Numerical Results
 7.1 Datasets
-We evaluate our methods on three multiple-choice QA datasets: CommonsenseQA (CQA) (Talmor et al., 2019), and the Social Sciences (SS) as well as STEM components from MMLU (Hendrycks et al., 2021).
+We evaluate our methods on three multiple-choice QA datasets: CommonsenseQA (CQA) Talmor et al.
+(2019), and the Social Sciences (SS) as well as STEM components from MMLU Hendrycks et al.
+(2021).
+The ensemble is trained on the CQA training dataset but tested on three different validation datasets: the (in-domain) CQA validation dataset and the (out-of-domain) MMLU STEM and MMLU SS validation datasets.
 Detailed information about the training and validation (test) splits is available in Appendix A.
 7.2 LoRA Ensembles
-We train LoRA ensembles using the pre-trained 7 billion parameter Mistral-7b model (Jiang et al., 2023).
-The LoRA parametrization was implemented using PEFT (Mangrulkar et al., 2022).
-Our ensembles contain 5 members $(M=5)$, but we also investigate larger ensembles in our ablation studies.
-Adhering to the method of Hu et al.
-(Hu et al., 2021), we applied the adapter $\Delta W=\alpha BA$ exclusively to the query and value matrices within the self-attention modules of Mistral-7b, with $\alpha$ set to 32.
-The rank was set to $r=8$, which gives $3,407,872$ trainable parameters.
-The adapter matrices $B$ are initialized to zero, while the entries of $A$ are initialized using a Kaiming-uniform distribution (He et al., 2015).
+We train LoRA ensembles using the pre-trained 7-billion-parameter Mistral-7B model Jiang et al.
+(2023).
+The LoRA parametrization is implemented using PEFT Mangrulkar et al.
+(2022).
+We evaluate ensembles with sizes $M\in\{1,5,10,20\}$ in our experiments and ablation studies.
+Following the method of Hu et al.
+Hu et al.
+(2021), we apply the adapter $\Delta W=\alpha BA$ exclusively to the query and value matrices within the self-attention modules of Mistral-7B, with $\alpha=32$.
+The adapter rank is set to $r=8$, resulting in $3{,}407{,}872$ trainable parameters.
+The adapter matrices $B$ are initialized to zero, and the entries of $A$ are initialized using a Kaiming-uniform distribution He et al.
+(2015).
 The ensemble members are trained using the prescribed regularisation from Section 5.3, corresponding to using the pretrained Mistral-7b model as a prior.
-We use the Adam optimizer, training for 10 epochs with learning step sizes of $5\cdot 10^{-6}$ and $1.25\cdot 10^{-6}$, and batch sizes of 8 and 2 for the CQA and MMLU datasets, respectively.
-These learning rates are smaller compared to those in (Wang et al., 2023) to better refine the learning curve before the models become over-fitted.
-The token size limit is fixed at 128 for CQA and 512 for MMLU datasets.
-We performed an ablation study on L2 regularization $\lambda/2=0.01,0.1,1,10$, finding optimal performance at $\lambda/2=1$ and $10$ for the CQA and MMLU datasets, respectively.
+We use the Adam optimizer, training for 10 epochs with learning step sizes of $5\cdot 10^{-6}$ and batch sizes of 8.
+This learning rate is smaller compared to those in Wang et al.
+(2023) to better refine the learning curve before the models become overfitted.
+The token size limit is fixed at 128 for CQA and 512 for MMLU.
+We performed an ablation study on L2 regularization $\lambda/2=0.01,0.1,1,10$, finding optimal performance at $\lambda/2=1$.
 The training focuses solely on the single token output representing the answer, with all other token outputs being masked.
@@
 This suggests that the model reliably comprehends the format of multiple-choice questions and answers, consistently producing appropriate tokens.
-Figure 1: Performance of the LoRA ensembles trained and evaluated on either CQA, MMLU STEM, or MMLU SS dataset.
-The metrics include accuracy, log likelihood loss, and expected calibration error (ECE).
-The number of ensemble members is either $M=5$ or $M=1$.
-For a LoRA ensemble with $M=1$ the average results over 5 distinct realizations are shown.
 7.3 Performance metrics
 In Fig.
-1, we illustrate the performance metrics for LoRA ensembles with both $M=1$ and $M=5$ members.
-The ensembles are trained on multiple-choice QAs from CQA (first column), MMLU STEM (second column), and MMLU Social Studies datasets (third column).
-The ensemble members start over-fitting after a couple of epochs as can be seen by the validation losses in the second row of panels, but interestingly without considerable drop in accuracy in the first row of panels.
+1, we present the performance metrics for LoRA ensembles with ensemble sizes $M=1$, $5$, $10$, and $20$.
+The ensembles are trained on multiple-choice questions from the CQA training dataset and evaluated on three validation datasets: CQA (first column), MMLU STEM (second column), and MMLU Social Studies (third column).
+As shown in Fig.
+1, the loss curves for ensemble sizes $M=5$, $10$, and $20$ converge closely, indicating that small ensembles ($M=5$) provide posterior approximations comparable to those of larger ensembles ($M=20$).
+This convergence is particularly interesting and practically relevant, as it suggests that LoRA ensembles can achieve accurate uncertainty estimates without significantly increasing inference costs.
+Based on this observation, we use $M=5$ ensembles for uncertainty quantification in the following section.
+Figure 1: Performance of LoRA ensembles trained on CQA and evaluated on the CQA, MMLU STEM, and MMLU SS datasets.
+Metrics (computed over the ensemble mean distribution) include accuracy, negative log-likelihood (NLL) loss, and expected calibration error (ECE).
+We show ensemble sizes $M\in\{1,5,10,20\}$; for $M=1,5,10$, results are averaged over 20, 4, and 2 independent realizations, respectively, while $M=20$ uses a single run.
+Metric values at epoch 1 (underfitting), epoch 3 (optimal), and epoch 6 (overfitting) are highlighted.
+The ensemble members start over-fitting after a couple of epochs as can be seen by the validation loss in the second row of panels, but interestingly without considerable drop in accuracy in the first row of panels.
 As will be seen in more detail in Section 7.4, this signals that the ensemble members become more overconfident on predictions that are wrong, corroborated by the increasing expected calibration error with epoch seen in the bottom row of Fig.
 1.
-This is consistent with the recent reports that fine-tuned LLMs often exhibit overconfidence (Jiang et al., 2021; Lin et al., 2022; Xiao et al., 2022b; He et al., 2023; Tian et al., 2023; OpenAI et al., 2023).
-For fine-tuning on CQA seen in the first column, the overfitting is significantly reduced for ensembles compared to individual members, as can be seen by the gap between the loss curves for $M=1$ and $M=5$ for later epochs.
+This is consistent with the recent reports that fine-tuned LLMs often exhibit overconfidence Jiang et al.
+(2021); Lin et al.
+(2022); Xiao et al.
+(2022b); He et al.
+(2023); Tian et al.
+(2023); OpenAI et al.
+(2023).
+For fine-tuning on CQA, shown in the first column, overfitting is significantly reduced for ensembles compared to individual models, as evidenced by the gap between the loss curves for $M=1$ and those for $M=5,10,20$ at later epochs.
+This suggests that ensemble members produce more confident yet diverse predictions.
 This means that ensemble members output more confident, but different predictions.
-This is a common feature of Bayesian posteriors (Blundell et al., 2015; Zhang et al., 2020; Kristiadi et al., 2020; Ober & Aitchison, 2021; Fortuin et al., 2022; Aitchison et al., 2021; Yang et al., 2024).
-This signals high epistemic uncertainty in this regime, indicating that the validation set is perceived as out-of-domain.
-There is no significant difference between single member and ensemble metrics for models trained on the MMLU datasets.
-We attribute this to the fact that the MMLU training datasets are small, see Appendix A, facilitating fast overfitting before any significant generalization is attained.
-Figure 2: AUROC computed for the CQA, MMLU STEM, and MMLU Social Studies multiple-choice QA datasets.
-In each subpanel, one dataset is used for fine-tuning (T) on the training subset and two datasets for the AUROC evaluation (V) on the validation subset.
-The LoRA ensemble size is $M=5$.
-7.4 Uncertainty metrics visualized by AUROC
-Area under receiver operator curve (AUROC) is commonly used as a metric for out-of-domain detection (Hendrycks & Gimpel, 2016), but can also be used more generally to measure how well a model can separate two datasets.
-By evaluating AUROC using uncertainty measures, we can gain understanding of which of the datasets is more consistently perceived by the model as more confusing (using entropy) or as more out-of-domain (using mutual information).
+This is a common feature of Bayesian posteriors Blundell et al.
+(2015); Zhang et al.
+(2020); Kristiadi et al.
+(2020); Ober & Aitchison (2021); Fortuin et al.
+(2022); Aitchison et al.
+(2021); Yang et al.
+(2024).
+This signals high epistemic uncertainty in this regime, indicating that the validation set is perceived as “unknown” to the model.
+7.4 Dynamics of uncertainty metrics
+We analyze data-resolved dynamics of uncertainty measures using two-dimensional histogram plots of mutual information and entropy Linander et al.
+(2023).
+This approach allows us to effectively disentangle different sources of uncertainty, providing a detailed view of how the model perceives data across the “known-certain,” “known-uncertain,” and “unknown” regimes during fine-tuning.
+By visualizing the relationship between mutual information and entropy, we can track how the model’s knowledge evolves, identify patterns in data complexity, and assess the model’s ability to adapt to fed information.
+This method not only highlights the dynamics of learning, but also helps identify limitations of the model, offering actionable insights for further improvement.
 In Fig.
-2, we show the AUROC results comparing in- and out-of-domain validation datasets.
-The ensembles were fine-tuned on the training subset of the CQA, MMLU STEM, and MMLU Social Studies datasets.
-For the AUROC evaluation, we use the validation subset of these datasets as in-domain samples, and the validation subset of other datasets as out-of-domain.
-By analyzing the AUROC curves for the ensemble trained on CQA in the first column of Fig.
-2, it is evident that the model consistently perceives samples from CQA as more in-domain compared to those from MMLU datasets, as indicated by an AUROC greater than 0.5.
-This aligns with expectations.
-However, the AUROC is not close to 1, suggesting that a significant portion of the MMLU dataset is perceived similarly to CQA in terms of entropy and mutual information, with MMLU Social Sciences appearing closer than MMLU STEM as can be seen by the lower AUROC values.
-A somewhat similar pattern is observed for the ensemble trained on MMLU Social Sciences.
-For the ensemble trained on MMLU STEM in the second column, a peculiar behavior is observed, where the AUROC is below 0.5.
-We interpret this as follows: MMLU STEM poses greater complexity for the LLM architecture, hence, even with training, it struggles to identify the right features for correct answers.
-Conversely, the ensemble retains robust knowledge about features pertinent to MMLU SS and CQA.
-A mutual information score below 0.5 indicates that the ensemble perceives the validation set of MMLU STEM as more out-of-domain compared to the other datasets, even though the model was fine-tuned on the training subset of MMLU STEM.
-7.5 Dynamics of uncertainty metrics
-To analyze the trends observed through the AUROC curves in more detail, we also produce data-resolved dynamics of uncertainty measures by looking at two-dimensional histogram plots of mutual information and entropy(Linander et al., 2023).
-In Fig.
-3, we show the epoch evolution of uncertainty measures for a LoRA ensemble that was fine-tuned and evaluated on the CQA dataset.
+2, we show the epoch evolution of uncertainty measures for a LoRA ensemble that was fine-tuned and evaluated on the CQA dataset.
 The mutual information (MI) on the vertical axis, and the predictive entropy (Entropy) on the horisontal axis, are binned over every sample of the validation set of CQA and the number of samples in each bin is indicated by the colorbar.
 We distinguish between incorrect and correct predictions to enhance interpretability.
-The ensemble attains minimal validation loss at epoch 3, as shown in Fig.
+The ensemble attains optimal validation loss at epoch 3, as shown in Fig.
 1.
 Epochs 1 and 6 exhibit underfitting and overfitting, respectively.
-The true posterior $p(\theta|D)$ should yield nearly zero mutual information when evaluated on the validation dataset, which should be in-domain by design.
-Samples with high mutual information indicate that parts of the target domain are perceived as out-of-domain by the model, showing the most important regions to consider for improving the training dataset targetting a better posterior approximation.
-We observe a moderate spread in entropy on the horizontal axis, indicating that the model predicts a broad distribution even though the sample is considered in-domain through their small mutual information.
-This spread can be attributed to the questions where the model does not have enough context or expressive power for a distinct answer.
+The true posterior $p(\theta|D)$ should yield nearly zero mutual information when evaluated on the validation dataset, which should belong to “known” data by design.
+However, in each of the underfitted, optimal, and overfitted regimes, there is a fraction of validation samples that exhibit a moderate spread in mutual information.
+Samples with high mutual information indicate that parts of the target domain are perceived as “unknown” by the ensemble, highlighting regions where expanding or refining the training data could most improve performance.
+By examining the mean and median of the entropic metrics in Fig.
+2, we can observe overall patterns in the validation data perception by the ensemble.
+Interestingly, the ensemble becomes progressively more confident on correct QAs, reaching its highest confidence once the ensemble overfits, where most correct QAs are predicted with almost zero uncertainty.
+Although overall performance (as measured by loss shown in Fig.
+1) degrades at epoch 6 compared to epoch 3, the model has actually learned the bulk of the validation dataset more effectively.
+The loss increase arises because the ensemble has also become more confident about incorrect QAs, increasing the cross-entropy.
+However, almost all of these incorrect predictions are marked by high mutual information, placing them in the “unknown” regime.
+Since these points can be identified by their high mutual information, they can be filtered out at inference time or used to refine the training procedure (e.g., by including more instances of these data).
+This result is noteworthy because it shows that the ensemble continues to improve even after reaching its minimum loss at epoch 3.
+By epoch 6, it performs better on most of the dataset, although its increased confidence in the remaining poorly performing data inflates the overall loss.
+Nonetheless, these problematic data points remain mostly distinguished by high epistemic uncertainty.
 Fig.
 3 shows the uncertainty histograms for MMLU STEM and MMLU SS, calculated using an ensemble trained on CQA.
-Starting with MMLU STEM in the first and second row of panels, we draw two conclusions.
-First, out of the wrongly classified samples for the optimal epoch 3 shown in the second row, second column, there is a large fraction with high predictive entropy and low mutual information compared to the more moderate spread of predictive entropy in the corresponding panel for MMLU SS in the last row, second column.
-This affirms conclusions from Section 7.4 regarding the relatively higher complexity of STEM compared to SS.
-The ensemble trained on CQA perceives the STEM samples equally in-domain as the SS samples, as evident by the similar mutual information, but predict a broader distribution indicating a lack of knowledge.
-Secondly, in the over-fitting regime corresponding to the last column, the mutual information of the wrongly classified samples of the second row increase more than the correctly classified samples in the first row.
-This indicates that the ensemble forgets the samples with high predictive entropy quicker, i.e.
-it forgets the confusing samples faster.
-Figure 3: Histograms of predictive entropy and mutual information for a LoRA ensemble trained and evaluated on the CQA dataset.
+Starting with MMLU STEM in the first and second row of panels, we draw two key insights.
+First, among the wrongly classified samples for the optimal epoch 3 (second row, second column), a large fraction falls into the “known-uncertain“ regime, characterized by high predictive entropy but low mutual information.
+This is in contrast to MMLU SS, which exhibits a more moderate spread of predictive entropy in the corresponding panel (last row, second column).
+This observation highlights the relatively higher complexity of STEM compared to SS datasets.
+The ensemble perceives the STEM samples as equally “known” as SS samples (similar mutual information) but assigns them broader predictive distributions, indicating greater “uncertain” characteristics due to a lack of precise knowledge.
+Figure 2: In-distribution evolution of uncertainty during fine-tuning.
+Histograms of predictive entropy and mutual information for a LoRA ensemble trained and evaluated on the CQA dataset.
 The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1.
-This figure illustrates the evolution of uncertainty measures for in-domain data across training epochs, differentiated by correct and incorrect predictions.
+This figure illustrates the evolution of uncertainty measures for in-domain data across training epochs (columns), differentiated by correct (top row) and incorrect (bottom row) predictions.
 We also depict the corresponding mean (red) and median (green) entropy and mutual information values.
-Figure 4: Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA and evaluated on MMLU STEM and MMLU SS datasets.
+Color represents the number count in each bin.
+Figure 3: Out-of-distribution evolution of uncertainty during fine-tuning.
+Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA and evaluated on MMLU STEM (top rows) and MMLU SS (bottom rows) datasets.
 The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1.
-This panels illustrates the evolution (left to right) of uncertainty measures for out-of-training-domain data across training epochs, differentiated by correct and incorrect predictions.
+These panels illustrates the evolution of uncertainty measures for out-of-training-domain data across training epochs (columns left to right), differentiated by correct (first and third row) and incorrect (second and forth row) predictions.
 We also depict the corresponding mean (red) and median (green) entropy and mutual information values.
-Continuing with MMLU SS in the third and fourth row, we observe a decrease in predictive entropy in the over-fitting regime, indicating that the ensemble does find features giving confident but wrong predictions.
-Interestingly, the evolution of entropic quantities shows a quantitative difference: For MMLU STEM, there is a significant density of high-entropy, low-mutual-information samples at epoch 6, whereas for MMLU SS, the data with high entropy is more dispersed in terms of mutual information.
-This suggests that the fine-tuned model perceives confusing samples from MMLU STEM and MMLU SS differently, forgetting the confusing samples from MMLU SS more rapidly.
+Color represents the number count in each bin.
+Figure 4: Late overfitting regime (epoch 10).
+Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA (left column) and evaluated on MMLU SS (middle column) and MMLU STEM (right column) datasets.
+The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1.
+There is still a significant fraction of samples (lower right panel) in MMLU STEM with low mutual information and high total uncertainty.
+Second, continuing with MMLU SS in the third and fourth rows of Fig.
+3, we observe a significant decrease in predictive entropy as the number of epochs increases, indicating that the ensemble learns features that enable it to classify samples with greater confidence.
+This behavior is similar to CQA in Fig.
+2, but the key difference is that, in this case, the data is from an out-of-domain dataset.
+This demonstrates that the ensemble continues to improve on the majority of MMLU SS samples, despite not being specifically trained on them, suggesting a similarity between the CQA and MMLU SS domains.
+Interestingly, the evolution of entropic quantities reveals a key difference between the datasets.
+For MMLU STEM, there is a notable density of “known-uncertain” samples (high entropy, low mutual information) (see Fig.
+3, second row), whereas for MMLU SS, high-entropy samples are more evenly distributed in terms of mutual information.
+The density of “known-uncertain” MMLU STEM samples largely persists despite overfitting (see Fig.
+4), suggesting that the fine-tuned model ensemble perceives confusing samples from MMLU STEM as being more consistently “known-uncertain,” whereas confusing samples from MMLU SS are forgotten more rapidly, transitioning into the “unknown” regime.
 8 Conclusions
 We derived a principled, and practically efficient method for posterior approximation of fine-tuned LLMs using ensembles of low-rank adaptation models.
-By analysing the evolution of entropic uncertainty measures during fine-tuning, we identified signals of dataset complexity and signs of architecture limitations.
-The methods presented are generally applicable to different fine-tuning scenarios, and provide a systematic treatment of prediction uncertainty for fine-tuned LLMs.
-9 Acknowledgements
-HL was supported by a grant from the KAW Foundation.
-The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at C3SE partially funded by the Swedish Research Council through grant agreements no.
-2022-06725 and no.
-2018-05973.
-OB would like to thank Ihor Balabanov for valuable discussions.
+Using the posterior approximation, we derived uncertainty quantification using entropic metrics, and followed their evolution during fine-tuning.
+This enabled us to draw a number of conclusions regarding the balance of retained knowledge and domain specific adaptation.
+First, the low-rank ensemble shows retained accuracy while also decreasing both predictive uncertainty and mutual information in the overfitting regime.
+At the same time, incorrect predictions are increasingly correlated with higher mutual information, opening up for inference time filtering and active learning.
+Second, during fine-tuning towards CQA, there is a distinct difference between the change in uncertainty for MMLU STEM and MMLU social science.
+For the STEM dataset, there is a large fraction of samples with low mutual information and high predictive entropy, hinting at in-domain but uncertain predictions.
+This can be connected to the lower prior accuracy of MMLU STEM compared to MMLU SS and CQA, indicating low representative power of the model in this domain.
+These STEM samples exhibit increased robustness, through broad yet consistent predictive distributions across ensemble members, to a large extent surviving overfitting.
+As fine-tuning continues to be an important tool in shaping model output Zhang et al.
+(2023b), it is also increasingly important to understand how the knowledge of the pre-trained model changes as the model is adapted to a target domain.
+The methods presented here are generally applicable to different fine-tuning scenarios, and provide a systematic treatment of prediction uncertainty for fine-tuned LLMs.
+9 Acknowledgments
+The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at C3SE partially funded by the Swedish Research Council through grant agreements 2022-06725 and 2018-05973.
+10 Impact Statement
+This paper advances the field of Machine Learning by enhancing the reliability and understanding of large language models (LLMs) through uncertainty quantification using LoRA ensembles.
+Our work contributes to the understanding and trustworthiness of LLM predictions, a step forward in AI reliability and applicability.
 Appendix A Datasets
 A.1 Size
-For MMLU STEM and Social Studies datasets, we merge the development and validation sets to create the training set and use the original test split as the validation set, as in (Wang et al., 2023).
+For MMLU STEM and Social Studies datasets, we use the original test split as the validation set, as in Wang et al.
+(2023).
 For CQA dataset we use the standard training and validation splits.
-The topics in the MMLU dataset are categorized into STEM, Social Studies, Humanities, and Other, following the classification defined in (Hendrycks et al., 2021).
+The topics in the MMLU dataset are categorized into STEM, Social Studies, Humanities, and Other, following the classification defined in Hendrycks et al.
+(2021).
 Dataset Train Size Validation Size
@@
 The datasets have been preprocessed into a single, consistently formatted prompt question-answer. Below, we present examples of these questions for each of the considered dataset.
+{mdframed}
+[backgroundcolor=gray!20, linecolor=black, linewidth=1pt, roundcorner=10pt]
 Q: A revolving door is convenient for two direction travel, but it also serves as a security measure at a what?
-Answer Choices: (a) bank (b) library (c) department store (d) mall (e) new york A: (a).
+Answer Choices:
+(a) bank
+(b) library
+(c) department store
+(d) mall
+(e) new york
+A: (a).
 CommonsenseQA
+{mdframed}
+[backgroundcolor=gray!20, linecolor=black, linewidth=1pt, roundcorner=10pt]
 Q: Which one of the following is the most appropriate definition of a 99% confidence interval?
-Answer Choices: (a) 99% of the time in repeated samples, the interval would contain the true value of the parameter (b) 99% of the time in repeated samples, the interval would contain the estimated value of the parameter (c) 99% of the time in repeated samples, the null hypothesis will be rejected (d) 99% of the time in repeated samples, the null hypothesis will not be rejected when it was false A: (a).
+Answer Choices:
+(a) 99% of the time in repeated samples, the interval would contain the true value of the parameter
+(b) 99% of the time in repeated samples, the interval would contain the estimated value of the parameter
+(c) 99% of the time in repeated samples, the null hypothesis will be rejected
+(d) 99% of the time in repeated samples, the null hypothesis will not be rejected when it was false
+A: (a).
 MMLU SS
-Q: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field. Answer Choices: (a) 0 (b) 1 (c) 2 (d) 3 A: (b). MMLU STEM
+{mdframed}
+[backgroundcolor=gray!20, linecolor=black, linewidth=1pt, roundcorner=10pt]
+Q: Find all c in Z_3 such that Z_3[x]/(x^2 + c) is a field.
+Answer Choices:
+(a) 0
+(b) 1
+(c) 2
+(d) 3
+A: (b).
+MMLU STEM